Rosa is on the school track team and ran the following distances during the past six days 5.6, 4.4, 3.6, 4.2, 5.3, 3.7 If she wantes to run at least a daily average of 4.7 miles over a 7-day period, what is the minimum number of miles she has to run on the seventh day?


My answer was 1.4 but the book has a different answer.

Found 3 solutions by sudhanshu_kmr, solver91311, AlgeGirl:
Answer by sudhanshu_kmr(1152)


Sum of 5.6, 4.4, 3.6, 4.2, 5.3, 3.7 is 26.8

4.7 * 7 = 32.9

minimum number of miles required for 7th day = 32.9 - 26.8 = 6.1


Answer by solver91311(24713)



The challenge for me on this one is trying to figure out the process you used to get to an answer of 1.4, and I confess to being utterly stumped. Except to say that you might have just reached around behind yourself and pulled a number out of your backside -- which would be one possible explaination of why your answer smells bad.

Consider how you go about calculating the average. You add up all of your data elements and then divide by the number of data elements. In order to get an average of 4.7 over 7 days, the sum of the seven data elements must be 4.7 times 7 which is 32.9. So whatever the sum of your first six data points happens to be, subtracted from the required sum for 7 data points, namely 32.9, must be the required number of miles to be run on the last day to achieve the desired average.

AH HA! I went back and found out what you did. You took the sum of the first six data points, 26.8, and then multiplied the required average, 4.7, by 6 to get 28.2, and then took the difference, 1.4. Close but no cigar. What you calculated was the average daily shortfall from a 4.7 average over the first 6 days. In order to achieve a 7 day average of 4.7, on the 7th day you would need to do the 4.7 PLUS that average daily shortfall of 1.4 for a total of 6.1. One way to get there, but the way I described in the previous paragraph is much easier and much less prone to error.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by AlgeGirl(1)